You — Please update your info
For the next fiscal​ year, you forecast net income of $ 50000 and ending assets of $ 500000. Your​ firm's payout ratio is 10.0 %. Your beginning​ stockholders' equity is $ 300000 and your beginning total liabilities are $ 120000. Your​ non-debt liabilities such as accounts payable are forecasted to increase by $ 10000. Assume your beginning debt is $ 100000. What amount of equity and what amount of debt would you need to issue to cover the net new financing in order to keep your​ debt-equity ratio​ constant?
1
Expert's answer
2018-09-26T09:54:08-0400
As Total assets = Total liability + Stockholders' Equity, so the beginning balance is 300000 + 120000 = 420000 and the current debt-equity ratio is: 120000/300000 = 0.4 or 40%.
Ending assets will be 500000, total liabilities will increase by 10000 and will be 130000, equity will increase by 50000*(1 - 0.1) = 45000 and will be 345000, so total equity and liabilities will be 345000 + 130000 = 475000, so, we need additional 25000 to make a balance.
If x is new debt and y is new equity, then to keep your debt-equity ratio constant:
x + y = 25000, x = 25000 - y,
(130000 + x)/(345000 + y) = 0.4,
(155000 - y)/(345000 + y) = 0.4,
0.4*(345000 + y) = 155000 - y,
1.4y = 155000 - 138000,
y = 17000/1.4 = 12143 - increase in equity needed.
x = 25000 - 12143 = 12857 - increase in debt needed